OpenStudy (lilah96):

If (f + g)(x) = 3x2 + 2x – 1 and g(x) = 2x – 2, what is f(x)? 3x2 + 1 3x2 – 1 3x2 + 4 3x2 – 4 x2 – 2

OpenStudy (awesomedude2078):

can you explain a bit

OpenStudy (visarika23):

(f+g)(x) is simply = f(x)+g(x) so f(x)=(f+g)(x)-g(x)

OpenStudy (lilah96):

do I just put the numbers in the f and x spots

OpenStudy (visarika23):

yup (3x^2+2x-1)-(2x-2)

OpenStudy (lilah96):

thank you

OpenStudy (visarika23):

wlcm:)